QUESTION 18 A metal object at 15.00 °C is heated by gaining 5.81 kJ of heat from the environment. The final temperature is 18NEED ANSWERS ASAP PLZ
Answer #1

Answer :- (E) Heat capacity of metal = 1.75 kJ/oC

Solution:-

for solving this question first we have to know the relation between heat and heat capacity then we have to use all the values

so as we know that

Heat = heat capacity * temperature change

or in the mathematical expression

Q = C * \Delta T..............(1)

it is given that

Q= 5.81 kJ

\DeltaT = T final - T initial = 18.32 - 15 = 3.32oC

put these values in equation (1)

C= 5.81 / 3.32

C = 1.75 kJ/oC

so the correct option is (E) 1.75 kJ/oC

please rate my answer